Give the domain of the function, f(x)= xsquare-5x-24 over xsquare-64
Answer by oscargut(686)

f(x)=(x^2-5x-24)/(x^2-64) the domain is D= x real numbers except the roots of x^2-64=0
so the domain is D = R-{8,-8} (Real numbers except 8 and -8)